OpenStudy (happykiddo):

he magnitude of displacement a and b are 3 m and 4 m respectively. Considering various orientation of a and b, what is i) maximum magnitude for c and ii) the minimum possible magnitude?

OpenStudy (happykiddo):

This is vectors and I just cant understand how...vector c minimum=a-b=3-4=1. How is it not -1?????

OpenStudy (happykiddo):

Never mind I figured it thanks to nobody who helped. This website is turning into rubbish. It used to be helpful before though : (